OpenStudy (anonymous):

what are 13 ways to make 100 out of 7 and 11?

OpenStudy (anonymous):

7 + 7 + 11 = 25 25 + 25 + 25 + 25 = 100 4(7 + 7 + 11) = 100 4(7 + 7 + 11) + 7 - 7 = 100 4(7 + 7 + 11) + 7 - 7 + 7 - 7 = 100 4(7 + 7 + 11) + 7 - 7 + 7 - 7 + 7 - 7 = 100 4(7 + 7 + 11) + 7 - 7 + 7 - 7 + 7 - 7 + 7 - 7 = 100 4(7 + 7 + 11) + 7 - 7 + 7 - 7 + 7 - 7 + 7 - 7 + 7 - 7 = 100 4(7 + 7 + 11) + 7 - 7 + 7 - 7 + 7 - 7 + 7 - 7 + 7 - 7 + 7 - 7 = 100 4(7 + 7 + 11) + 7 - 7 + 7 - 7 + 7 - 7 + 7 - 7 + 7 - 7 + 7 - 7 + 7 - 7 = 100 4(7 + 7 + 11) + 7 - 7 + 7 - 7 + 7 - 7 + 7 - 7 + 7 - 7 + 7 - 7 + 7 - 7 + 7 - 7 = 100 4(7 + 7 + 11) + 7 - 7 + 7 - 7 + 7 - 7 + 7 - 7 + 7 - 7 + 7 - 7 + 7 - 7 + 7 - 7 + 7 - 7 = 100 4(7 + 7 + 11) + 7 - 7 + 7 - 7 + 7 - 7 + 7 - 7 + 7 - 7 + 7 - 7 + 7 - 7 + 7 - 7 + 7 - 7 + 7 - 7 = 100 4(7 + 7 + 11) + 7 - 7 + 7 - 7 + 7 - 7 + 7 - 7 + 7 - 7 + 7 - 7 + 7 - 7 + 7 - 7 + 7 - 7 + 7 - 7 + 7 - 7 = 100 4(7 + 7 + 11) + 7 - 7 + 7 - 7 + 7 - 7 + 7 - 7 + 7 - 7 + 7 - 7 + 7 - 7 + 7 - 7 + 7 - 7 + 7 - 7 + 7 - 7 + 7 - 7 = 100